November 03, 2009 - NMDB phone conference

Short reports for period October 16 - October 31, 2009

CAU

  • participation in the October telecon
  • participation statistics for training course
  • set up new webserver for www.nmdb.eu in Kiel (Drupal 6.14)
  • set up replacement server for db04
  • added BM35 and GB1 barometer, WS-2307 Weather station and Heinzinger NCE 3000-10 HV power supply to Registration Systems library
  • webpage to display GCR parameters calculated by IZMIRAN: www.nmdb.eu/gcr/

OULU

  • Participation in October telecon
  • Relocation of the project website to Kiel
  • Improvement of website features (translations, email notifications) (presently underway)
  • Relocation of some Online Tools from Oulu to Kiel
  • Translation of PO documentation (brochure) into Finnish

IZMIRAN

  • Participation in October telecon
  • Work on the reports on Deliverables WP5.1, WP5.4.2, WP5.5. Loading these reports on the nmdb website.
  • Finetuning of the webpages of the Internet projects (improving the Tables, adding some explanations, comments and so on)
  • Additional work on the uploading data from NMDB for usage them in the Applications (due to some changes in the NEST program).
  • Repeating upload the Reports and some other documents to the NMDB website because of this site shifting.
  • Preparing a documentation for Library on WP5.

NKUA

  • Participation in October telecon
  • Preparation of a paper on NMDB for submission at the Proceedings vol of the 9th Helaset Conference which will be published by the Astronomical Society of the Pacific (deadline November 15, 2009)
  • Keeping the Athens mirror up to date with security updates and bug fixes on packages that we use on NMDB (phpmyadmin interface).
  • Changed the networks' parameters on the server due to changes of university's local network.
  • Working on correcting our station data and towards the improvement of the data correction on the database.
  • Resolved a problem on our second counter that appeared on October 27 13:00. The problem was fixed on October 28 15:00
  • Made discussions about the Registration systems Library and working on increasing its uselessness and attractiveness
  • Improving the interface of the GLE alert system with database
  • Working on improvements of the GLE alert program
  • Working on the connection of the GLE alert to NM Bangle software
  • Working on the translation of PO documentation (send by O. Kryakunova) into Greek.
  • Preparations on the final meeting (3-5 December 2009) which will be held at Athens

Alcalá

  • Participation en October telecon
  • Software for data retrieval
  • Software NMDB data insertion
  • FPGA changes to allow serial port multiplexing

OBSPARIS

  • Organisation of, participation in NMDB telecon on Oct 16
  • NMDB training course: some administrative work
  • Work on training material (reading of received documents), minor work on PO
  • Nest tool: add the monthly sunspot index and smoothed sunspot index as an overplotting option. Transfer of nest to www.nmdb.eu and correction of minor compatibility issues. UTC bug correction.
  • LK & NF attended the EURADOS meeting, whose members are interested in the NMDB achievements.
  • A local mysql database is under development. It will facilitate the transfer of historical data and future real time data to NMDB.

UBERN

  • Participation in NMDB telecon on Oct 16
  • Deliverable 5.3.3: Follow-up development of webpage to publish the radiation dose rates and the ionization rates, starting to write the documentation for deliverable 5.3.3
  • Deliverable 6.5: Request to IZMIRAN and NKUA to document their part in determination of the parameters that describes the GCR and SCR flux near Earth. IZMIRAN and NKUA will send their contribution until 13 November 2009 to UBERN. UBERN has started to write their part.
  • Deliverable 6.11: Translation of PO webpage to German together with CAU.

Aragats

Summary:Aragats group is progresses according to the schedule. Before the starting of Winter operation power supplies and tubes were checked. One failed tube at Aragats was changed to a new one. The failed tube at Nor Ambered was checked; power was changed and now it is operating properly. 2 new powerful servers (HP and IBM cell) are put into operation and data migration software are under preparation now. The downloading of the neither revised data for Nor Amberd Neutron Monitor to NMDB is finished. All tables are now available from NMDB.
  • (WP3) Electronics and tubes we checked. One tube at Aragats was changed, one channel at Nor Amberd was repaired. Now all 36 channels of Aragats and Nor Amberd neutron monitors are operational and are sending data to NMDB.
  • (WP4 ) New data servers were commissioned and installed at CRD. The LINUX (UBUNTU) operational systems were installed. The migration tool for transferring data files from old dataserver to main Data Storiage Server (DSS) is under preparation.
  • (WP5) The examples of the library of the NMDB applications documents are prepared.
  • (WP6) The design of the NMDB educational site was tuned according to the request of the WP6 package leader.

Kosice

  • Participation in NMDB telecon on Oct 16
  • WP5. Implementation of discrete FT into provisional delivery 5.6. Delivery 5.6 is checked for correctness, working version at http://www.nmd.eu/~igor/variation
  • WP6. Part of PO website translated to Slovak. Continuing .

ESOI-TAU

  • Participation in NMDB telecon on Oct 16
  • Modernization software for data transfer from ESOI observatory to NMDB on the case of short time electric blackout, caused by atmospheric discharge and change electric supply on diesel-generator. This update increased access to ESOI data on 24 hours.
  • Modernization of software with transfer real time minute data to site ESOI (both graphics and ASCII data.
  • Part of the WP6 matter is translated to hebrew. Remained part in progress.
  • In the last days we meet with problem of electricity supply on ESOI observatory caused by electricity blackout of Mount Hermon caused by first winter urican with great atmospheric electric discharge. This storm weather situation may continue up to 03/11/2009. For saving of activity of the observatory in extremal weather condition we plan in nearest weeks to increase time work of reserve diesel generator up to 4-5 days.

Almaty

  • Participation in NMDB telecon on Oct 16
  • added Cockroft-Walton type HV power supply to Registration Systems library
  • Translation of PO documentation into Kazakh
  • Calculation of hourly and daily Alma-Ata NM data for 1973-2009 (WP5.6)

UNIROMA3

  • Partecipation in the telecon on 16 October.
  • Correction of the new tool realised for sending data to NMDB. Improving of algorithm test for taking into account all requested recommendations from NMDB.
  • Analysis of the IDL 6.0 version to compute the cosmic ray diurnal variations. We found that this version is fully compatible with the opensource GDL program and therefore can solve the licence problem.